# Opponent Shaping for Antibody Development

### v4

@@ Abstract
Anti-viral therapies are typically designed to target the current strains of a virus. Game theoretically, this corresponds to a short-sighted, or myopic, response. However, therapy-induced selective pressures act on viral antigens to drive the emergence of mutated strains, against which initial therapies have reduced efficacy. Building on a computational model of binding between antibodies and viral antigens (the Absolut! framework), we design and implement a genetic simulation of such viral evolutionary escape. Crucially, this allows our antibody optimisation algorithm to consider and influence the entire escape curve of the virus, i.e. to guide (or “shape”) the viral evolution. This is inspired by opponent shaping which, in general-sum learning, accounts for the adaptation of the co-player rather than playing a myopic best response. Hence we call the optimised antibodies shapers. Within our simulations, we demonstrate that our shapers target both current and simulated future viral variants, outperforming the antibodies chosen in a myopic way. Furthermore, we show that shapers exert specific evolutionary pressure on the virus compared to myopic antibodies. Altogether, shapers modify the evolutionary trajectories of viral strains and minimise the viral escape compared to their myopic counterparts. While this is a simplified model, we hope that our proposed paradigm will enable the discovery of better long-lived vaccines and antibody therapies in the future, enabled by rapid advancements in the capabilities of simulation tools. Our code is available at https://github.com/olakalisz/antibody-shapers.
@@ 1 Introduction
Building on these principles, our work focuses specifically on antibody design, framing the interaction between antibodies and surface viral antigens as a two-player zero-sum game. In this game, the antibody’s payoff is primarily determined by its binding strength to the virus, whilst the virus has the opposite payoff. Although our framework can use any binding model (or indeed any sequence-to-function map), in this paper we use our GPU-accelerated implementation of the Absolut! framework [11] to approximate the binding strength of protein-protein interactions. We use this game to model viral escape - the process via which mutations in a virus allow it to successfully evade a host’s immune system [12]. In the language of game theory, this is the virus adapting to the current antibody by repeatedly finding a local best response, decreasing the virus’s binding strength to the antibody. Using this viral escape model we optimise antibodies to be effective over the entire process of viral evolution, resulting in designed antibodies which we refer to as shapers. This is in contrast to only optimising for binding to the initial virus, which results in myopic antibodies.
Importantly, our simulations reveal that myopic antibodies are suboptimal when considering a longer time horizon of viral evolution. Shaper antibodies not only outperform their myopic counterparts in long-term efficacy but also exert a shaping influence on viral evolutionary trajectories, effectively constraining the virus’s ability to escape antibody binding. Our study also explores the trade-offs between the effectiveness of shaper antibodies and the computational resources required for their optimisation, providing insights into the scalability and potential for practical deployment of our method. Finally, we present an analysis of the key features that distinguish between shaper and myopic antibodies.
In summary, this work introduces a novel antibody design philosophy leveraging opponent-shaping principles to improve the long-term efficacy of antibodies against evolving viruses. While our results provide a promising proof of concept, they are based on simplified models of binding and viral escape, limiting direct real-world application. However, as more sophisticated methods for modelling protein interactions and predicting viral evolution emerge, our framework has the potential to significantly impact future antiviral therapy design.
@@ 2 Results
Figure 1: Computational Implementation of Opponent Shaping for Antibody Design. a An overview of our implementation. In the Antibody Optimisation Loop (i.e. the outer loop), we optimise the antibody to perform well against the current and future virus variants and influence the viral evolution. We approximate the future variants through the Simulated Viral Escape via Evolution (i.e. the inner loop) where the viral antigens are evolved against the current antibody over a given horizon length. b The payoffs of the antibody and virus. Red arrows indicate binding interactions that players aim to minimise, while green arrows represent those they aim to maximise. The antibody optimises for binding to the virus while avoiding its anti-target. In this zero-sum game, the antibody’s optimisation indirectly counters the virus’s binding to its target. c Efficient JAX [13] implementation of the binding calculation uses binding poses generated by Absolut! [11] and the Miyazawa-Jernigan energy potential matrix [14].
@@ 2.1.2 Simulated Viral Escape via Evolution
Viral Escape is the process of a virus adapting to a host’s immune system, eventually becoming fully resistant. From a game theory perspective, this is the process of the virus “learning to play the game”. To simulate the space of possible viral escape trajectories against a chosen antibody, we employ an evolutionary algorithm. This approach approximates the process of viral evolution, where mutations that increase fitness (in our case, the virus’s payoff $R_{v}$) are more likely to persist, propagate, and become an ancestor of future generations.
We simulate this process of viral escape for a specified horizon of $H$ steps, with each step representing a single generation of viral evolution. This is the target duration of therapy efficacy. For each independent viral trajectory, we use the initial virus strain $v$ to generate a population of $P$ virus variants through random mutations. These are point mutations in the amino acid sequence of the viral antigen and they are introduced according to a predefined mutation rate. Each variant’s fitness is evaluated based on its payoff $R_{v}$ against the fixed antibody $a$. Importantly, this process captures the myopic nature of viral evolution: the virus optimises its fitness only with respect to the current antibody, unable to anticipate future human interventions or antibody changes. The next generation of the viruses is then created by probabilistically selecting the variants based on their fitness, with fitter variants more likely to be chosen for replication and subsequent mutation. This process repeats for $H$ generations, creating one trajectory of viral escape, see the Simulated Viral Escape via Evolution in Figure 1a and Methods section 4.2 for more details.
We use the antigen protein from the Dengue Virus as the virus for all our experiments, specifically, the structure with Protein Data Bank (PDB) code 2R29 [21, 22]. We discretise this structure using the Absolut! framework to generate the docking poses, which we then use to calculate the binding strength $B$ of any antibody-virus pair using our JAX implementation. In the viral escape step, we mutate only the amino acid sequence of the dengue envelope antigen, which is composed of 97 amino acids, and do not consider other components of Dengue Virus. Importantly, while we alter the amino acid sequence, we assume the structure of the antigen does not significantly change for the calculation of binding strength.
@@ 2.1.3 Optimisation of Antibody Shapers
Our antibody optimisation process aims to generate antibodies that can prevent viral escape. In particular, we desire antibodies which maintain their performance even once the virus has had the opportunity to evolve in response to them. For this purpose, we define the antibody fitness function $F^{H}_{v}(a)$ as the antibody’s performance over multiple potential viral escape trajectories of virus $v$:
@@
Specifically, for each of the $\eta$ trajectories, we simulate the viral escape process for $H$ steps and calculate the average antibody payoff $R_{a}$ over these steps, where $v^{\prime}$ represents the possible mutations of the original virus $v$. The final fitness is the average over these $\eta$ simulations, i.e. $\eta$ Monte Carlo roll-outs. The fitness function $F^{H}_{v}(a)$ captures the average antibody’s long-term effectiveness against evolving viral populations.
We then employ an evolutionary algorithm on the antibodies themselves to optimise this fitness function. In particular, optimisation proceeds iteratively for $N$ steps where we select the antibody with the highest fitness at each step and generate a new population of $P_{a}$ antibodies by introducing single mutations to the selected antibody. This process of evaluation, selection, and mutation continues, gradually improving the antibody’s robustness against, and ability to limit, viral escape, see Antibody Optimisation Loop in Figure 1a and Methods section 4.3 for more details.
We refer to the resulting optimised antibodies as shapers, or more specifically as $H=5$ shapers or $H=100$ shapers, reflecting the horizon length chosen for the viral escape simulations during their optimisation. Myopic antibodies are antibodies where horizon length $H=0$ and thus $F^{0}_{v}(a)=R_{a}(v,a)$. As we will discuss in more detail, the choice of $H$ significantly impacts the antibody’s long-term effectiveness at the cost of computational resources required for optimisation [^1] A single step of antibody optimisation requires simulating $P_{a}\times\eta$ independent viral escapes, each for horizon $H$ steps. However, we can parallelise over both $P_{a}$ and $\eta$., making horizon $H$ a vital parameter to consider.
@@ 2.2 Antibody Shapers Minimise Viral Escape
Figure 2: Optimised antibody shapers outperform myopic antibodies. a Distribution of antibody shapers (in orange) which were optimised with horizon $H=100$ vs. myopic antibodies distribution (in blue). We highlight the top 10% shapers with respect to $F^{100}_{v}(a)$ in red and the top 10% myopic antibodies with respect to $R_{a}(v,a)$ in green. The x-axis is the myopic antibody fitness $R_{a}(v,a)$ and the y-axis is the escape averaged antibody fitness over 100 steps of viral escape. Higher values on both axes indicate better performance. b Viral escape curves for different steps of the antibody optimisation process for antibody shapers optimised with horizon 100 (solid lines) and myopic antibodies (dashed lines). The lighter lines indicate early antibody optimisation steps and the darker lines show the later steps. The x-axis shows the evolutionary steps of viral escape. The y-axis represents the virus fitness/payoff $R_{v}(v,a)$, where higher values indicate better virus fitness (and lower values denote better antibody performance. c, d Antibody optimisation learning curves for a varying horizon length. The x-axis shows the antibody optimisation steps (c) and the number of samples from the binding simulator (d). The y-axis represents the escape averaged antibody fitness over 100 steps of viral escape $F^{100}_{v}(a)$. Error bars correspond to the standard error. Higher values indicate better performance.
Our results demonstrate a clear advantage for antibody shapers in the escape-averaged objective $F^{100}_{v}(a)$. The mean of the shapers distribution significantly exceeds that of the myopic distribution, as evident from the marginal density plot in Figure 2a. Notably, none of the myopic antibodies outperforms the top 10% of shapers in this long-term objective. However, there is a trade-off between short-term and long-term optimisation. While shapers do better on the escape-averaged objective, they underperform compared to myopic antibodies in optimising the immediate payoff $R_{a}(v,a)$ against the initial, non-mutated virus $v$.
@@ 2.2.2 Antibody Shapers with Varying Horizons
However, the number of antibody optimisation steps does not accurately reflect the computational or experimental cost of optimisation. Each simulation of viral escape requires a number of binding samples that increases linearly with the horizon length $H$. Yet, shorter horizon antibodies optimise an objective that diverges further from our “true” antibody objective $F^{100}_{v}(a)$. Due to this trade-off, we observe that the optimal training horizon varies depends on the available computational budget.
To illustrate this trade-off, we conduct an additional experiment shown in Figure 2d. Here, instead of fixing the number of optimisation steps $N$, we constrain the total number of binding samples — queries to our binding strength simulator used to evaluate all antibody and virus payoffs throughout the optimisation process — to be constant across different horizons. This approach provides a performance comparison that accounts for the computational resources necessary across varying horizon lengths. Interestingly, $H=20$ shapers perform strongly - nearly matching the performance of those optimised with horizon $H=100$ for a given number of antibody optimisation steps, and far exceeding it when accounting for the differing computational cost. This suggests that using a cheaper, shorter-horizon, proxy for the true antibody objective $F^{100}_{v}(a)$ can yield substantial benefits.
@@ 2.3.1 Attack is the Best Defence
Figure 3: Effect of myopic and shaper antibodies on viruses induced by other antibodies. We optimise $80$ different antibodies $a_{H}$ for each horizon (Myopic, $H=5,H=10,H=20,H=100$), these are represented by the y-axis. We simulate the viral escape to each of these antibodies for $100$ steps and we group the escaped viruses $v_{H}$ by the horizon $H$ of the antibody that induced them, these escape viruses are represented by the x-axis. In colour, we show the mean antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$ for each group of optimised antibodies $a_{H^{\prime}}$ against the final escape viral variant $v_{H}$ induced by other antibodies optimised with horizon $H$. Darker colours correspond to better antibody payoff.
To disentangle these strategies, we seperately evaluate the antibodies and the viruses that evolve in response to them. To do this, we compare the virus against other antibodies which did not influence the viral evolution. The intuition is that an antibody which is good at shaping (good attack), but less robust (poor defence), will induce viruses which other antibodies will perform well against. Specifically, we generate antibodies $a_{H}$ for each horizon $H$ and simulate viral escape against these antibodies for 100 steps, resulting in viruses $v_{H}$. For all pairs of horizons $(H,H^{\prime})$, we then cross-evaluate the antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$. Figure 3 presents the result of this analysis.
@@ 2.3.3 Influence of Antibody Shapers on Binding Poses
Figure 5a presents average pose matrices from multiple optimisation runs of both myopic antibodies and long-horizon shapers. We observe the following two trends. We observe two key trends. First, as viral escape steps increase (top row vs bottom row), the pose matrices become more “diffused”. This is expected, as the virus explores more “pose possibilities” through mutations during escape. Second, as the horizon of antibody optimisation increases, the poses also become more “diffused”. This is particularly interesting, as all antibodies have the same number of mutations regardless of the horizon, suggesting that this diffusion might relate to the increased robustness of shapers.
To further understand these pose dynamics, we aggregate the pose matrices along the antibody axis (Figure 5b) and the virus axis (Figure 5c). These figures show the change in interaction counts between viral escape steps 00 and $100$. Figure 5b shows that as the virus escapes it includes more of the antibody’s lowest binding amino acids (particularly K) in the pose. Notably, long-horizon shapers, especially $H=100$ shapers, are most effective at preventing this increase in K interactions. Furthermore, Figure 5c shows another viral escape strategy, where the virus removes its high-binding amino acids (I and M) from the pose. Again, $H=100$ shapers are most successful in counteracting this trend, although they cannot completely prevent it.

### v6

@@ Abstract
Anti-viral therapies are typically designed to target only the current strains of a virus. Game theoretically, this corresponds to a short-sighted, or myopic, response. However, therapy-induced selective pressures act on viruses to drive the emergence of mutated strains, against which initial therapies have reduced efficacy. Building on a computational model of binding between antibodies and viral antigens (the Absolut! framework), we design and implement a genetic simulation of viral evolutionary escape. Crucially, this allows our antibody optimisation algorithm to consider and influence the entire escape curve of the virus, i.e. to guide (or “shape”) the viral evolution. This is inspired by opponent shaping which, in general-sum learning, accounts for the adaptation of the co-player rather than playing a myopic best response. Hence we call the optimised antibodies shapers. Within our simulations, we demonstrate that our shapers target both current and simulated future viral variants, outperforming the antibodies chosen in a myopic way. Furthermore, we show that shapers exert specific evolutionary pressure on the virus compared to myopic antibodies. Altogether, shapers modify the evolutionary trajectories of viral strains and minimise the viral escape compared to their myopic counterparts. While this is a simplified model, we hope that our proposed paradigm will facilitate the discovery of better long-lived vaccines and antibody therapies in the future, enabled by rapid advancements in the capabilities of simulation tools. Our code is available at https://github.com/olakalisz/antibody-shapers.
@@ 1 Introduction
Building on these principles, our work focuses specifically on antibody design, framing the interaction between antibodies and surface viral antigens as a two-player zero-sum game. In this game, the antibody’s payoff is primarily determined by its binding strength to the virus, whilst the virus has the opposite payoff. Although our framework can use any binding model (or indeed any sequence-to-function map), in this work we use our GPU-accelerated implementation of the Absolut! framework [11] to approximate the binding strength of protein-protein interactions. We use this game to model viral escape - the process via which mutations in a virus allow it to successfully evade a host’s immune system [12]. In the language of game theory, this is the virus adapting to the current antibody by repeatedly finding an approximate best response, decreasing the virus’s binding strength to the antibody. Using this viral escape model we optimise antibodies to be effective over the entire process of viral evolution, we refer to these as shapers. This is in contrast to only optimising for binding to the initial virus, which results in myopic antibodies.
Importantly, our simulations reveal that myopic antibodies are suboptimal when considering a longer time horizon of viral escape. Shaper antibodies not only outperform their myopic counterparts in long-term efficacy but also exert a shaping influence on viral evolutionary trajectories, effectively constraining the virus’s ability to escape antibody binding. Our study also explores the trade-offs between the effectiveness of shaper antibodies and the computational resources required for their optimisation, providing insights into the scalability and potential for practical deployment of our method. Finally, we present an analysis of the key features that distinguish shapers from myopic antibodies.
In summary, this work introduces a novel antibody design philosophy and implementation framework, leveraging opponent-shaping principles to improve the long-term efficacy of antibodies against evolving viruses. While our results provide a promising proof of concept, they are based on simplified models of binding and viral escape, limiting direct real-world application. However, as more sophisticated methods for modelling protein interactions and predicting viral evolution emerge, our framework has the potential to significantly impact future antiviral therapy design.
@@ 2 Results
Figure 1: Computational Implementation of Opponent Shaping for Antibody Design. a An overview of our implementation. In the Antibody Optimisation Loop (i.e. the outer loop), we optimise the antibody to perform well against current and future virus variants; thus influencing the viral evolution. We approximate the future variants through our Simulated Viral Escape via Evolution (i.e. the inner loop) where the viruses evolve to escape from the current antibody over a given horizon length. b The payoffs of the antibody and virus. Red arrows indicate binding interactions that players aim to minimise, while green arrows represent those they aim to maximise. The antibody optimises for binding to the virus while avoiding its anti-target. In this zero-sum game, the antibody’s optimisation indirectly counters the virus’s binding to its target, see Equation 1. c Efficient JAX [13] implementation of the binding calculation uses binding poses generated by Absolut! [11] and the Miyazawa-Jernigan energy potential matrix [14].
@@ 2.1.2 Simulated Viral Escape via Evolution
Viral Escape is the process of a virus adapting to evade the host’s immune system [12]. From a game theory perspective, this is the process of the virus “learning to play the game”. To simulate the space of possible viral escape trajectories against a chosen antibody, we employ an evolutionary algorithm. This approach approximates the process of viral evolution, where mutations that increase fitness (in our case, the virus’s payoff $R_{v}$) are more likely to persist, propagate, and become an ancestor of future generations.
We simulate this process of viral escape for a specified horizon of $H$ steps, with each step representing a single generation of viral evolution. This is the target duration of therapy efficacy. For each independent viral trajectory, we use the initial virus strain $v$ to generate a population of $P$ virus variants through random mutations. These are point mutations in the amino acid sequence of the viral antigen and they are introduced according to a predefined mutation rate. Each variant’s fitness is evaluated based on its payoff $R_{v}$ against the fixed antibody $a$. Importantly, this process captures the myopic nature of viral evolution: the virus optimises its fitness only with respect to the current antibody, unable to anticipate future human interventions or antibody changes. The next generation of viruses is then created by probabilistically selecting the variants based on their fitness, with fitter variants more likely to be chosen for replication and subsequent mutation. This process repeats for $H$ generations, creating one trajectory of viral escape, see the Simulated Viral Escape via Evolution in Figure 1a and Methods Section 4.2 for more details.
As an example implementation of our method we use the antigen protein from the Dengue Virus for all our experiments, specifically, the structure with Protein Data Bank (PDB) code 2R29 [21, 22]. We discretise this structure using the Absolut! framework to generate the docking poses, which we then use to calculate the binding strength $B$ of any antibody-virus pair using our JAX implementation. In the viral escape step, we mutate only the amino acid sequence of the dengue envelope antigen, which is composed of 97 amino acids, and do not consider other components of Dengue Virus. Importantly, while we alter the amino acid sequence, we assume the structure of the antigen does not significantly change for the calculation of binding strength.
@@ 2.1.3 Optimisation of Antibody Shapers
Our antibody optimisation process aims to generate antibodies that can prevent viral escape. In particular, we want antibodies which maintain their performance even once the virus has had the opportunity to evolve in response to them. For this purpose, we define the antibody fitness function $F^{H}_{v}(a)$ as the antibody’s average performance over multiple simulated viral escape trajectories of virus $v$:
@@
Specifically, for each of the $\eta$ trajectories, we simulate the viral escape process for $H$ steps and calculate the average antibody payoff $R_{a}$ over these steps, where $v^{\prime}$ represents the possible mutated variants of the original virus $v$. The final fitness is the average over these $\eta$ simulations, i.e. $\eta$ Monte Carlo roll-outs. The fitness function $F^{H}_{v}(a)$ approximates the antibody’s long-term effectiveness against evolving viral populations.
We then employ an evolutionary algorithm on the antibodies themselves to optimise this fitness function. In particular, starting from random antibodies at step 00, optimisation proceeds iteratively for $N$ steps where we select the antibody with the highest fitness at each step and generate a new population of $P_{a}$ antibodies by introducing single mutations to the selected antibody. This process of evaluation, selection, and mutation continues, gradually improving the antibody’s robustness against, and ability to limit, viral escape. See Antibody Optimisation Loop in Figure 1a and Methods Section 4.3 for more details.
We refer to the resulting optimised antibodies as shapers, or more specifically as $H=5$ shapers or $H=100$ shapers, reflecting the horizon length of the viral escape simulations used during their optimisation. Myopic antibodies are antibodies where horizon length $H=0$ and thus $F^{0}_{v}(a)=R_{a}(v,a)$. As we will discuss in more detail, the choice of $H$ significantly impacts the antibody’s long-term effectiveness at the cost of computational resources required for optimisation [^1] A single step of antibody optimisation requires simulating $P_{a}\times\eta$ independent viral escapes, each for horizon $H$ steps. However, we can parallelise over both $P_{a}$ and $\eta$., making horizon $H$ a vital parameter to consider.
@@ 2.2 Antibody Shapers Minimise Viral Escape
Figure 2: Optimised antibody shapers outperform myopic antibodies. a Distribution of antibody shapers (in orange) which were optimised with horizon $H=100$ vs. myopic antibodies distribution (in blue). We highlight the top 10% shapers with respect to $F^{100}_{v}(a)$ in red and the top 10% myopic antibodies with respect to $R_{a}(v,a)$ in green. The x-axis is the myopic antibody fitness $R_{a}(v,a)$ and the y-axis is the escape averaged antibody fitness over 100 steps of viral escape. Higher values on both axes indicate better performance. b Viral escape curves for different steps of the antibody optimisation process for antibody shapers optimised with horizon 100 (solid lines) and myopic antibodies (dashed lines). The lighter lines indicate early antibody optimisation steps and the darker lines show the later steps. The x-axis shows the evolutionary steps of viral escape. The y-axis represents the virus fitness/payoff $R_{v}(v,a)$, where higher values indicate better virus fitness (and lower values denote better antibody performance). c, d Antibody optimisation learning curves for a varying horizon length. The x-axis shows the antibody optimisation steps (c) or the number of samples from the binding simulator (d). The y-axis represents the escape averaged antibody fitness over 100 steps of viral escape $F^{100}_{v}(a)$. Error bars correspond to the standard error. Higher values indicate better performance.
Our results demonstrate a clear advantage of shapers in the escape-averaged objective $F^{100}_{v}(a)$. The mean of the shapers distribution significantly exceeds that of the myopic distribution, as evident from the marginal density plot in Figure 2a. Notably, none of the myopic antibodies outperforms the top 10% of shapers in this long-term objective. However, there is a trade-off between short-term and long-term optimisation. While shapers do better on the escape-averaged objective, they underperform compared to myopic antibodies in optimising the immediate payoff $R_{a}(v,a)$ against the initial, non-mutated virus $v$.
@@ 2.2.2 Antibody Shapers with Varying Horizons
However, the number of antibody optimisation steps does not accurately reflect the computational or experimental cost of optimisation. Each simulation of viral escape requires a number of binding samples that increases linearly with the horizon length $H$. Yet, shorter horizon antibodies optimise an objective that diverges further from our “true” antibody objective $F^{100}_{v}(a)$. Due to this trade-off, we observe that the optimal training horizon varies depending on the available computational budget.
To illustrate this trade-off, we conduct an additional experiment shown in Figure 2d. Here, instead of fixing the number of optimisation steps $N$, we constrain the total number of binding samples - queries to our binding strength simulator used to evaluate all antibody and virus payoffs throughout the optimisation process - to be constant across different horizons. This approach provides a performance comparison that accounts for the computational resources necessary across varying horizon lengths. Interestingly, $H=20$ shapers perform strongly, nearly matching the performance of those optimised with horizon $H=100$ for a given number of antibody optimisation steps, and far exceeding it when accounting for the differing computational cost. This suggests that using a cheaper, shorter-horizon, proxy for the true antibody objective $F^{100}_{v}(a)$ can yield substantial benefits.
@@ 2.3.1 Attack is the Best Defence
Figure 3: Effect of myopic and shaper antibodies on viruses induced by other antibodies. We optimise $80$ different antibodies $a_{H}$ for each horizon (Myopic, $H=5,H=10,H=20,H=100$), these are represented by the y-axis. We simulate the viral escape to each of these antibodies for $100$ steps and we group the escape viruses $v_{H}$ by the horizon $H$ of the antibody that induced them, these escape viruses are represented by the x-axis. In colour, we show the mean antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$ for each group of optimised antibodies $a_{H^{\prime}}$ against the final escape viral variant $v_{H}$ induced by other antibodies optimised with horizon $H$. Darker colours correspond to better antibody payoff.
To disentangle these strategies, we seperately evaluate the antibodies and the viruses that evolve in response to them. To do this, we compare the viruses against other antibodies which did not influence the viral evolution. The intuition is that an antibody which is good at shaping (good attack), but less robust (poor defence), will induce viruses which other antibodies will perform well against. Specifically, we generate antibodies $a_{H}$ for each horizon $H$ and simulate viral escape against these antibodies for 100 steps, resulting in viruses $v_{H}$. For all pairs of horizons $(H,H^{\prime})$, we then cross-evaluate the antibody payoff $R_{a}(v_{H},a_{H^{\prime}})$. Figure 3 presents the result of this analysis.
@@ 2.3.3 Influence of Antibody Shapers on Binding Poses
Figure 5a presents average pose matrices from multiple optimisation runs of both myopic antibodies and long-horizon shapers. We observe two key trends. First, as viral escape steps increase (top row vs bottom row), the pose matrices become more “diffused”. This is expected, as the virus explores more “pose possibilities” through mutations during escape. Second, as the horizon of antibody optimisation increases, the poses also become more “diffused”. This is particularly interesting, as all antibodies have the same number of mutations regardless of the horizon, suggesting that this diffusion might relate to the increased robustness of shapers.
To further understand these pose dynamics, we aggregate the pose matrices along the antibody axis (Figure 5b) and the virus axis (Figure 5c). These figures show the change in interaction counts between viral escape steps 00 and $100$. Figure 5b shows that as the virus escapes it includes more of the antibody’s lowest binding amino acids (particularly K, Lysine) in the pose. Notably, long-horizon shapers, especially $H=100$ shapers, are most effective at preventing this increase in K (Lysine) interactions. Furthermore, Figure 5c shows another viral escape strategy, where the virus removes its high-binding amino acids I (Isoleucine) and M (Methionine) from the pose. Again, $H=100$ shapers are most successful in counteracting this trend, although they cannot completely prevent it.